W varies directly as x and y and inversely as v square and w =1200 when x =4, y=9 and v=6.find w when p =6 and s =4.

Answer by Alan3354(69443)

W varies directly as x and y and inversely as v square
W = k*x*y/y^2 where k = a constant.
----------
and w =1200 when x =4, y=9 and v=6.
1200 = k*4*9/36
k = 1200
----
find w when p =6 and s =4.
----------
p and s are not defined.